OpenStudy (anonymous):

For which x values does the graph of y=4^x cross the x-axis? a. none b. negative values of x c. x values less that 1 d. positive values of x

OpenStudy (jamesj):

If the graph of this function crossed the x axis, then for some x, 4^x = 0. Is there such a number x?

OpenStudy (anonymous):

0?, so none?

OpenStudy (jamesj):

Well, 4^0 = 1. In fact, there is no such x. Hence y = 4^x is never zero. Hence ...

OpenStudy (anonymous):

if by none you mean answer A, you are correct

OpenStudy (anonymous):

yes, thank you for the explanation

OpenStudy (anonymous):

the range of the exponential function is y>0, i.e., strictly greater than zero